can someone help me with these problems:
(9x+2)=5[11-2(2-x)+3]

Answer by stanbon(48546)

(9x+2)=5[11-2(2-x)+3]
9x+2 = 5[11-4+2x+3]
9x+2 = 5[10+2x]
9x+2 = 50+10x
x=-48
============
Cheers,
Stan H.